(1 point) For a random variable X, suppose that E[X] = 2 and Var(X) = 3. Then (a) E[(5 + x)2) = (b) Var(2 + 6X) =
Answer #1

Solution :

E(X) = 2

V(X) = 3

V(X) = E(X2) - [E(X)]

E(X) = V(X) + [E(X)] = 3+22 = 3+4 = 7

a)

E [(5+ x)2) = E(25+X? + 10X) = 25+ E(X) + 10E(X)

E [(5+ X) = 25+ 7+10*2 = 52

b)

V(2 +6X) = 62V(X) = 36 * 3 = 108